On May 1, 2021, Hecala Mining entered into an agreement with the state of New Mexico to obtain the rights to operate a mineral mine in New Mexico for $9.5 million. Additional costs and purchases included the following (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.): Development costs in preparing the mine $ 2,700,000 Mining equipment 122,500 Construction of various structures on site 45,000 After the minerals are removed from the mine, the equipment will be sold for an estimated residual value of $10,000. The structures will be torn down. Geologists estimate that 750,000 tons of ore can be extracted from the mine. After the ore is removed the land will revert back to the state of New Mexico. The contract with the state requires Hecala to restore the land to its original condition after mining operations are completed in approximately four years. Management has provided the following possible outflows for the restoration costs: Cash Outflow Probability $ 550,000 30% 650,000 30% 750,000 40% Hecala’s credit-adjusted risk-free interest rate is 8%. During 2021, Hecala extracted 115,000 tons of ore from the mine. The company’s fiscal year ends on December 31.

During 2022, Hecala changed its estimate of the total amount of ore originally in the mine from 750,000 to 950,000 tons. Calculate the depletion of the mine and depreciation of the mining facilities and equipment for 2022 assuming Hecala extracted 145,000 tons of ore in 2022. (Do not round your intermediate calculations. Round "Depreciation" and "Depletion" rates to 4 decimal places. Round your final answers to the nearest whole dollar.)

Show less

2022
Depletion
Depreciation of equipment
Depreciation of structures
Answer #1

Assuming that depletion of the mine will be done in 'quantity produced' method, because total amount of mineral extracted and extraction per year are available. Because of no other information is available on the question it is assuming that depreciation of equipment and structures are done in straight line method.

computation of depletion of mineral mine

Depletion base= Total cost for purchase the right + development cost + restoration cost

restoration cost will be expected value of possible outcomes. Therefore restoration cost

=( 550000*30% + (650000*30%) +(750000*40%)= $ 660000

therefore depletion base = 9500000+2700000+660000= $12860000

Depletion for 2021 =Depletion base * Quantity extracted during the year / Total quantity extracted

= 12860000* 115000/ 750000

( quantity extracted during 2021 = 115000 tons, total quantity of ore in the mine = 750000 tons)

= $ 1971866.6667

it is given that during the year 2022 the estimation of amount of the ore extracted changed from 750000 to 950000 tons

so updated depletion base for 2022 = $12860000- $ 1971866.6667 = $ 10888133.33

remaining amount of ore can be extracted after estimate change = total ore can be extracted - extracted in 2021

=950000-115000= 835000 tons

Depletion for 2022 = 10888133.33*145000/ 835000= 1890753.6920 ( ore extracted during 2022 = 145000 tons)

Depreciation expense of equipment

= total cost - salvage value / life of the asset = 122500-10000/4 = $ 28125

total cost of equipment = $122500 salvage value after minerals removed = 10000 , the contract period for the mining operation will be the life of the equipment that is 4 years

Depreciation expense of structures

total construction cost of structures = $ 45000 salvage value - nil

life will be same as of contract

therefore depreciation = 45000/4 $4 = $ 11250

2022
Depletion 1890753.6920
Depreciation of equipment 28125
Depreciation of structures 11250